for 2018, what is the projected operating profit given the following information for a doctors office?

the office is expecting 4,300 patient visits in 2019 and has estimated the cost function as follows:

y=$128,004 + $15.60x
where “y” is the total cost and “x” is number of visits. Budgets revenue is $57 per visit. This is a non profit entity and does not pay taxes.

1.) whats the operating profit for 2019?
2.) breakeven point, margin of safety in units and sales $?

please supply step by step instructions
Answer #1

1)Under cost function equation : $ 128004 is a fixed cost and $ 15.60 per visit is a variable cost that varies with number of visit

sales revenue (57*4300) 245100
less:variable cost (15.60*4300) (67080)
Contribution margin 178020
less:fixed cost (128004)
net operating income 50016

2)contribution per visit =selling price -variable cost

                           = 57 -15.60

                            = $ 41.40 per visit

contribution margin ratio =contribution /price

                                      = 41.40 / 57

                                       = 72.63158%

i)Breakeven point (units) =Fixed cost /contribution per visit

                                   = 128004 /41.40

                                   = 3092 visits

Breakeven point ($) = 128004/.7263158

                           = $ 176237

ii)Margin of safety (units) =Actual sales -Breakeven point sales

                               = 4300-3092

                               = 1208 visits

iii)Margin of safety ($) = Actual sales -Breakeven sales

                         = 245100-176237

                          =$ 68863